# Movable Signals with Dual-Polarized Fixed Intelligent Surfaces: Beyond Diagonal Reflection Matrices

### v1

@@ I Introduction
Reconfigurable intelligent surfaces (RISs) have been thoroughly investigated as a promising means to enable smart radio environments (SREs) by shaping and controlling the wireless channel [1, 2]. A RIS is a surface composed of reflecting elements with adjustable electromagnetic (EM) properties, thereby steering incident signals toward desired directions. Although RIS can offer interesting performance gains, their practical deployment remains challenging. Dynamic reconfiguration of the RIS requires dedicated control circuitry with non-negligible power consumption. An effective RIS reconfiguration relies on accurate channel knowledge, demanding additional training overhead. Moreover, since a RIS alters all incident EM signals, the resulting unpredictable interference from different operators raises concerns.
To circumvent these limitations, an alternative approach has been introduced: movable signals operating with fixed intelligent surface (FIS) [3]. Instead of reconfiguring the EM properties of the surface, this approach reconfigures the carrier frequency of the transmitted signal, while using a surface of uniformly spaced reflecting elements with fixed EM properties. Prior work has demonstrated that, in single-input single-output (SISO) systems, movable signals with FIS can enhance the channel strength by up to four times compared with RIS. Although this gain relies on a suitable spectrum availability (ideally, the available carrier frequencies should be distributed over a wide range), it remarkably does not require any reconfigurable hardware or additional channel estimation overhead.
@@ II-A Channel Model
The channels $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$ are modeled by assuming that the surface is a dual-polarized uniform linear array (ULA) in line-of-sight (LoS) with both the transmitter and receiver. The $N$ elements of the surface are numbered such that the first $N/2$ have vertical polarization and the last $N/2$ have horizontal polarization, as shown in Fig. 1. The effect of the element polarization is incorporated in $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$ by modeling them as
@@
The LoS fading channels $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$ are expressed as follows. We place the $N$ elements of the surface on the $x$-axis, centered in $x=0$, so that the $n$th element has $x$ coordinate $x_{n}=(n-(N+1)/2)d_{A}$, with $d_{A}$ being the spacing between the antenna elements. Denoting as $d_{R}$ the distance between the center of the surface and the receiver, and as $\theta_{R}\in[-\pi/2,\pi/2]$ the angle of the receiver with respect to the surface normal, the distance between the $n$th element and the receiver is $d_{R,n}=d_{R}-x_{n}\sin(\theta_{R})$ assuming far-field conditions. Hence, the entries of $\mathbf{g}_{R}$, given by $\mathbf{g}_{R}]_{n}=e^{-j2\pi d_{R,n}/\lambda}$, can be written as
@@
as a function of the wavelength $\lambda=c/f$, where $c$ the speed of light, and $f$ the frequency. Likewise, denoting as $d_{T}$ distance between the center of the surface and the transmitter, and as $\theta_{T}\in[-\pi/2,\pi/2]$ the angle of the transmitter, the entries of $\mathbf{g}_{T}$ write as
@@ II-B Problem Formulation
where $P_{T}=\mathbb{E}[|x|^{2}]$ denotes the transmitted power, assumed to be $P_{T}=1$ for simplicity. We want to compare two strategies for maximizing $P_{R}$.
@@ II-B1 RIS
With a RIS, the received power is maximized by reconfiguring $\boldsymbol{\Theta}$ depending on the channels $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$, which are assumed to be known [1]. Formally, the optimization problem is
@@ III-1 Diagonal RIS
$\bar{P}_{R}^{\mathrm{DRIS}}=\frac{\left(1+\chi\right)^{2}}{4}\left(N^{2}+4N\mathbb{E}\left[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right]+4\mathbb{E}[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2}]\right),$ (9)
where $\mathbf{g}_{R}\mathbf{g}_{T}$ can be approximated as a complex Gaussian with mean 0 and variance $N/2$, i.e., $\mathbf{g}_{R}\mathbf{g}_{T}\sim\mathcal{CN}(0,N/2)$, by the Central Limit Theorem. Hence, exploiting $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|]=\sqrt{\pi N/8}$ and $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|^{2}]=N/2$, we obtain
@@ III-3 Diagonal FIS
The optimality of this value can be verified by substituting $\lambda^{\star}$ into the expressions of $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$ in (2) and (3) and verifying that it leads to $|\mathbf{g}_{R}\mathbf{g}_{T}|=N/2$. Therefore, by reconfiguring the signal frequency as in (15) a D-FIS with reflection matrix $\boldsymbol{\Theta}=-\mathbf{I}_{N}$ can achieve a received power $P_{R}^{\mathrm{DFIS}}=(1+\chi)^{2}N^{2}$. Note that this value is four times higher than the received power achieved by a D-RIS $\bar{P}_{R}^{\mathrm{DRIS}}\approx(1+\chi)^{2}N^{2}/4$, and can be obtained without reconfiguring the EM properties of the surface. Movable signals can achieve higher performance than RIS since they control both terms $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$ and $\mathbf{h}_{R}\mathbf{h}_{T}$ in the channel expression, while RIS only control $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$.
@@ IV-4 Beyond-Diagonal FIS
which means we are considering a group-connected architecture with $N/2$ groups, each including 2 elements [11]. The groups are formed by two co-located elements having opposite polarization and all have a $2\times 2$ scattering matrix given by $0,-1]^{T},[-1,0]^{T}]$. Since an example of $2$-port network having scattering matrix $0,-1]^{T},[-1,0]^{T}]$ is a phase shifter with phase $\pi$ (see Fig. 2(a)), such a BD-FIS can be implemented by connecting each pair of co-located elements via a phase shifter with phase $\pi$ (see Fig. 2(b)). In this way, the signal incident on the vertically polarized elements is reflected from the horizontally polarized ones, and vice versa, and the BD-FIS behaves as a polarization converter. Note that (29) is a block diagonal matrix with $2\times 2$ blocks being $0,-1]^{T},[-1,0]^{T}]$ if its columns are permuted. With (29), the received power is
@@ V Gain over Diagonal RIS
Focusing on systems where the transmitter and receiver have opposite polarization, we report the gains $G^{\mathrm{BDRIS}}$, $G^{\mathrm{DFIS}}$, and $G^{\mathrm{BDFIS}}$ as a function of $\chi$ in Fig. 3. We make the following three observations. First, the gain of BD-RIS over D-RIS tends to infinity when $\chi=0$, decreases with $\chi$, and is $1$ when $\chi=1$. BD-RIS is therefore particularly beneficial for small values of $\chi$, as also noticed in [10]. Second, the gain of D-FIS over D-RIS is always $4$ and does not vary with $\chi$, as for uni-polarized systems [3]. Third, the gain of BD-FIS over D-RIS tends to infinity when $\chi=0$, decreases with $\chi$, and reaches $4$ when $\chi=1$. Therefore, BD-FIS with movable signals is an appealing strategy for achieving substantially higher performance than RIS in dual-polarized systems. Since its reconfigurability lies in the signal frequency, BD-FIS has a fixed reflection matrix, which is easy to implement and does not require control hardware.
@@ VI Conclusion
We have investigated the performance limits of SISO systems aided by dual-polarized intelligent surfaces. Specifically, we have compared two strategies: RIS, where the reflection properties of the surface are reconfigured, and movable signals with FIS, where the signal frequency is reconfigured while the surface has fixed reflection properties. We show that movable signals with FIS always outperform RIS, by achieving at least a four times higher received power. When the transmitter and receiver have opposite polarization, a FIS with a reflection matrix not limited to be diagonal, namely a BD-FIS, remarkably further increases the performance. Future research could explore BD-FIS architectures that are optimal for any polarization of the transmitting and receiving antennas, and the optimization of the frequency of movable signals under realistic spectrum constraints and channel models.

### v2

@@ I Introduction
Reconfigurable intelligent surfaces (RISs) have been thoroughly investigated as a promising means to enable smart radio environments (SREs) by shaping and controlling the wireless channel [1, 2]. A RIS is a surface composed of reflecting elements with adjustable electromagnetic (EM) properties, thereby steering incident signals toward desired directions. Although RIS can offer interesting performance gains, its practical deployment remains challenging. Dynamic reconfiguration of the RIS requires dedicated control circuitry with non-negligible power consumption. An effective RIS reconfiguration relies on accurate channel knowledge, demanding additional training overhead. Moreover, since a RIS alters all incident EM signals, the resulting unpredictable interference from different operators raises concerns.
To circumvent these limitations, an alternative approach has been introduced: movable signals operating with fixed intelligent surface (FIS) [3]. Instead of reconfiguring the EM properties of the surface, this approach reconfigures the carrier frequency of the transmitted signal, while using a surface of uniformly spaced reflecting elements with fixed EM properties. Prior work has demonstrated that, in single-input single-output (SISO) systems, movable signals with FIS can enhance the channel strength by up to four times compared with RIS [3]. Although this gain relies on a suitable spectrum availability (ideally, the available carrier frequencies should be distributed over a wide range), it remarkably does not require any reconfigurable hardware or high channel estimation overhead.
@@ II-A Channel Model
The channels $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$ are modeled by assuming that the surface is a dual-polarized uniform linear array (ULA) in line-of-sight (LoS) with both the transmitter and receiver. The $N$ elements of the surface are numbered such that the first $N/2$ have vertical polarization and the last $N/2$ have horizontal polarization, as shown in Fig. 1. The effect of the element polarization is captured in $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$ by modeling them as
@@
The LoS fading channels $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$ are expressed as follows. We place the $N$ elements of the surface on the $x$-axis, centered in $x=0$, so that the $n$th element has $x$ coordinate $x_{n}=(n-(N+1)/2)d_{A}$, with $d_{A}$ being the spacing between the antenna elements. Denoting as $d_{R}$ the distance between the center of the surface and the receiver, and as $\theta_{R}\in[-\pi/2,\pi/2]$ the angle of the receiver with respect to the surface normal, the distance between the $n$th element and the receiver is $d_{R,n}=d_{R}-x_{n}\sin(\theta_{R})$ assuming far-field. Hence, the entries of $\mathbf{g}_{R}$, given by $\mathbf{g}_{R}]_{n}=e^{-j2\pi d_{R,n}/\lambda}$, are
@@
as a function of the wavelength $\lambda=c/f$, where $c$ the speed of light, and $f$ the frequency. Likewise, denoting as $d_{T}$ the distance between the center of the surface and the transmitter, and as $\theta_{T}\in[-\pi/2,\pi/2]$ the angle of the transmitter, the entries of $\mathbf{g}_{T}$ are written as
@@ II-B Problem Formulation
where $P_{T}=\mathbb{E}[|x|^{2}]$ is the transmitted power, assumed $P_{T}=1$ for simplicity. We compare two strategies for maximizing $P_{R}$.
@@ II-B1 RIS
With a RIS, the received power is maximized by reconfiguring $\boldsymbol{\Theta}$ depending on the channels $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$, which are assumed to be known [1]. Formally, the problem is
@@ III-1 Diagonal RIS
Taking the expectation of (8) over $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$, whose entries are assumed to have independent phases uniformly distributed in $0,2\pi)$, the average received power $\bar{P}_{R}^{\mathrm{DRIS}}=\mathbb{E}[P_{R}^{\mathrm{DRIS}}]$ is
$\bar{P}_{R}^{\mathrm{DRIS}}=\frac{\left(1+\chi\right)^{2}}{4}\left(N^{2}+4N\mathbb{E}\left[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right]+4\mathbb{E}[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2}]\right).$ (9)
Furthermore, since $\mathbf{g}_{R}\mathbf{g}_{T}$ is the sum of $N/2$ independent and identically distributed (i.i.d.) random variables with mean 0 and variance 1, it can be approximated as a complex Gaussian with mean 0 and variance $N/2$, i.e., $\mathbf{g}_{R}\mathbf{g}_{T}\sim\mathcal{CN}(0,N/2)$, by the Central Limit Theorem. Hence, by exploiting $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|]=\sqrt{\pi N/8}$ and $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|^{2}]=N/2$, we obtain
@@ III-3 Diagonal FIS
The optimality of this value can be verified by substituting $\lambda^{\star}$ into the expressions of $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$ in (2) and (3) and verifying that it leads to $|\mathbf{g}_{R}\mathbf{g}_{T}|=N/2$. Therefore, by reconfiguring the signal frequency as in (15) a D-FIS with reflection matrix $\boldsymbol{\Theta}=-\mathbf{I}_{N}$ can achieve a received power $P_{R}^{\mathrm{DFIS}}=(1+\chi)^{2}N^{2}$. Note that this value is four times higher than the received power achieved by a D-RIS $\bar{P}_{R}^{\mathrm{DRIS}}\approx(1+\chi)^{2}N^{2}/4$, and can be obtained without reconfiguring the EM properties of the surface. Movable signals can achieve higher performance than RIS since they control both terms $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$ and $\mathbf{h}_{R}\mathbf{h}_{T}$ in the received power expression, while RIS only controls $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$.
@@ IV-4 Beyond-Diagonal FIS
which means we are considering a group-connected architecture with $N/2$ groups, each including 2 elements [11] (note that (29) is a block diagonal matrix with $2\times 2$ blocks being $0,-1]^{T},[-1,0]^{T}]$ if its columns are permuted). The groups are formed by two co-located elements having opposite polarization and all have a $2\times 2$ scattering matrix given by $0,-1]^{T},[-1,0]^{T}]$. Since an example of $2$-port network having scattering matrix $0,-1]^{T},[-1,0]^{T}]$ is a phase shifter with phase $\pi$ (Fig. 2(a)), such a BD-FIS can be implemented by connecting each pair of co-located elements via a phase shifter with phase $\pi$ (Fig. 2(b)). In this way, the signal incident on the vertically polarized elements is reflected from the horizontally polarized ones, and vice versa, and the BD-FIS is a polarization converter. With (29), the received power is
@@ V Gain over Diagonal RIS
Focusing on systems where the transmitter and receiver have opposite polarization, we report the gains $G^{\mathrm{BDRIS}}$, $G^{\mathrm{DFIS}}$, and $G^{\mathrm{BDFIS}}$ as a function of $\chi$ in Fig. 3. We make the following three observations. First, the gain of BD-RIS over D-RIS tends to infinity when $\chi=0$, decreases with $\chi$, and is $1$ when $\chi=1$. BD-RIS is therefore particularly beneficial for small values of $\chi$, as also explained in [10]. Second, the gain of D-FIS over D-RIS is always $4$ and does not vary with $\chi$, as for uni-polarized systems [3]. Third, the gain of BD-FIS over D-RIS tends to infinity when $\chi=0$, decreases with $\chi$, and reaches $4$ when $\chi=1$. Therefore, BD-FIS with movable signals is an appealing strategy for achieving substantially higher performance than RIS in dual-polarized systems. Since its reconfigurability lies in the signal frequency, BD-FIS has a fixed reflection matrix, which is easy to implement and does not require control hardware.
@@ VI Conclusion
We have investigated the performance limits of SISO systems aided by dual-polarized intelligent surfaces. Specifically, we have compared two strategies: RIS, where the reflection properties of the surface are reconfigured, and movable signals with FIS, where the signal frequency is reconfigured while the surface has fixed reflection properties. We show that movable signals with FIS always outperform RIS, by achieving at least a four times higher received power. When the transmitter and receiver have opposite polarization, a FIS with a reflection matrix not limited to be diagonal, namely a BD-FIS, remarkably further increases the performance. Future research could explore BD-FIS architectures that are optimal for any polarization of the transmitter and receiver, and the optimization of the frequency of movable signals under realistic spectrum constraints and channel models, e.g., in the near-field.